OpenStudy (anonymous):

Solve: 2sin(x) = 3^(1/2)tan(x) So my teacher says in some cases we can't cancel the common factor or something :P so is this applicable here?? does this simplify to cos(x) = (3^0.5)/2 ?? Cause if it does, then I don't need help :)

OpenStudy (anonymous):

and @hartnn I spoke to my teacher: he said it's a typo xD *fist-in-the-air* ACHIEVEMENT UNLOCKED. My teacher admitted defeat. HA. I'm so happy :D

OpenStudy (anonymous):

Is the right side, (3^1/2) times tangent, or is tangent also part of the exponent on the 3?

OpenStudy (anonymous):

nopee, its times tan, tan's not part of the index.

OpenStudy (anonymous):

ok I get ur question. We couldn't cancel here too. There is sin (x) =0

OpenStudy (anonymous):

Your solution looks correct then. Might want to still check your solutions, though, just in case.

OpenStudy (anonymous):

i got \[\pm 30, \pm 330\]

OpenStudy (anonymous):

^^ right, that would make 0 also a solution...

OpenStudy (anonymous):

sin(0)=tan(0)

OpenStudy (anonymous):

but i simplified it. Then x would not be 0.

OpenStudy (anonymous):

Put x=0 into the original equation and check.

OpenStudy (anonymous):

(and x=π . . .)

OpenStudy (anonymous):

in the original equation it does work.

OpenStudy (anonymous):

The original equation is what matters most. The simplified forms are equivalent for some of the solutions, and can give more information, but they sometimes take away information too. The original equation is always boss.

OpenStudy (anonymous):

\[2\sin(x)=\sqrt{3}\tan(x)\] \[2\sin(x)-\sqrt{3}\tan(x)=0\] \[2\sin(x)-\frac{ \sqrt{3}\sin(x) }{ \cos(x) }=0\] \[\sin(x)(2-\frac{ \sqrt{3} }{ \cos(x) })=0\]\[\sin(x)=0 ......and .....2-\frac{ \sqrt{3} }{ \cos(x) }=0\]Now solve for x.

OpenStudy (anonymous):

\[0,\pi/6,11\pi/6,\pi,2\pi\]are the answer for \[0<x<2\pi\]

OpenStudy (anonymous):

nopess, answer is for -360 - 360 inclusive.

OpenStudy (anonymous):

omg guys you know what, you've been really helpful, but i've got a few more questions to do, and I REALLY don't feel like changing it now. so i'll leave it as cos(x), and let the teacher mark me wrong :P BUT THANK YOU SO MUCHHH .

OpenStudy (anonymous):

A similar thing happens with rational equations. If you cancel common factors top-and-bottom, you can lose information about one of the discontinuities. e.g. \[\large y=\frac{x^2+3x+2}{x^2-4} \rightarrow y=\frac{(x+2)(x+1)}{(x+2)(x-2)} \] \[\large \rightarrow y=\frac{\cancel{(x+2)}(x+1)}{\cancel{(x+2)}(x-2)} \rightarrow y=\frac{(x+1)}{(x-2)}\] If you just go by the final equation, you'll miss the x=/=-2 hole in the graph.

OpenStudy (anonymous):

Just add the - ve sign for what I posted before and include them too. Thar will be the answers.